OpenStudy (anonymous):

express 4 8/9 as an improper fraction

OpenStudy (morganchavez):

44/9

OpenStudy (anonymous):

thank you so much got it right

OpenStudy (igreen):

To convert it into an improper fraction, multiply the denominator by the whole number, and then add it to the numerator. This will be your numerator value. You just keep your denominator the way it is. \(4 \dfrac{8}{9}\) Multiply 9 * 4: 9 * 4 = 36 Add 8: 36 + 8 = 44. So we have \(\dfrac{44}{9}\).

OpenStudy (igreen):

@morganchavez You're not supposed to give out direct answers without explanation.